OpenStudy (anonymous):

What is the simplified form of the expression below (6 + 7b) + 14

OpenStudy (raden):

just miss the bracket, and combine the similar term : (6 + 7b) + 14 = 6 + 14 + 7b = 20 + 7b

OpenStudy (anonymous):

thank you